OpenStudy (anonymous):

Which graph represents the function f(x) = x2 − 2x − 8?

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ y= x^2 − 2x − 8 }\) \(\normalsize\color{blue}{ y = x^2 − 2x \color{red}{ +1-9 } }\) \(\normalsize\color{blue}{ y = x^2 − 2x +1\color{red}{ -9 } }\) \(\normalsize\color{blue}{ y = (x-1)^2-9 }\) does this tell you anything ?